# Rethinking Visual Dependency in Long-Context Reasoning for Large Vision-Language Models

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) excel in cross-model tasks but experience performance declines in long-context reasoning due to overreliance on textual information and reduced visual dependency. In this study, we empirically analyze LVLMs in long-context reasoning, revealing that increased context length leads to a higher dependence on language at the expense of visual dependency. To address this issue, we propose a novel training-free context pruning method that selectively removes less critical textual information. Our approach enhances visual dependency and reduces textual noise, thereby improving LVLM performance in long-context reasoning. We validate our method by constructing a long-context dataset, demonstrating its effectiveness across various LVLMs. Moreover, further analysis confirms the robustness of different token pruning strategies and preliminary explores scaling laws between pruning rates and context length.
@@
Yucheng Zhou1, Zhi Rao2, Jun Wan2, Jianbing Shen1thanks:  Corresponding author. 1 SKL-IOTSC, CIS, University of Macau 2 Macau University of Science and Technology yucheng.zhou@connect.um.edu.mo, jianbingshen@um.edu.mo
@@ 1 Introduction
Recently, Large Language Models (LLMs) have significantly enhanced their performance in long-context reasoning through contextual position encoding and optimized attention strategies (Ding et al., 2024; Jin et al., 2024). These methods enable LLMs to retain and effectively utilize more information over long-context interactions, thereby maintaining contextual coherence and improving response accuracy. In contrast, LVLMs struggle to leverage visual information appropriately when processing long contexts. Consequently, it is essential for LVLMs to accurately integrate visual and textual information under long-context reasoning.
@@
To improve the visual dependency in long-context reasoning, we propose a novel training-free context pruning method for LVLMs, which strategically reduces the interference of less critical textual information while preserving visual information. Our approach leverages the model’s internal attention mechanisms to identify and eliminate the trivial textual tokens based on their aggregated attention weights across multiple attention heads. By systematically pruning these less significant tokens, our method effectively diminishes textual noise and encourages more attention to visual information, enhancing visual dependency under long-context reasoning. In addition, we provide a theoretical foundation demonstrating that context pruning leads to more concentrated and stable attention distributions, which in turn fosters a stronger visual dependency on visual inputs.
To comprehensively evaluate the performance of LVLMs in long-context reasoning, we construct a long-context dataset based on the SVIT dataset (Zhao et al., 2023). Experimental results demonstrate the effectiveness of our approach across various LVLMs. Additionally, we further verify the robustness of our method by comparing different token pruning strategies, as well as various pruning layers and rates. The characteristics of our pruning method are further analyzed through part-of-speech observations of the pruned tokens. Lastly, we conduct a preliminary exploration of the scaling laws about the relationship between token pruning rates and context length.
@@
  • We perform a theoretical analysis demonstrating that context pruning leads to more concentrated and stable attention distributions, which enhances visual dependency.
@@
2 Related Work
2.1 Large Vision-Language Models
Recent advancements in LVLMs have focused on enhancing cross-modal understanding by integrating visual and textual inputs for complex tasks (Wang et al., 2023; Chen et al., 2023a, b; Qu et al., 2024). For instance, LLaVA (Liu et al., 2023) leverages GPT-4 (OpenAI, 2023) to generate multimodal instruction-following data, improving performance in tasks like Science QA and multimodal reasoning. Expanding on this, models such as SVIT (Zhao et al., 2023) and LLaVA-NeXT (Liu et al., 2024a) utilize larger datasets and improve capabilities in OCR and abstract reasoning. Approaches like DeepSeek-VL (Lu et al., 2024) address real-world challenges by handling noisy and ambiguous inputs, while ViP-LLaVA (Cai et al., 2024) enhances robustness in processing diverse visual prompts. Additionally, some methods such as SURf’s self-refinement framework improve the accuracy of information retrieval (Sun et al., 2024), and VICL (Zhou et al., 2024) enables few-shot learning through inference on visual captioning. Zhang et al. (2024) extends to zero-shot video understanding using AnyRes, allowing models trained on images to generalize to videos, while Video-LLaMA2  (Cheng et al., 2024) enhances spatial-temporal modeling by integrating visual and audio inputs. These developments highlight LVLMs’ trends in scaling data, improving reasoning, and increasing robustness.
2.2 Long-Context Reasoning
Handling long contexts is a key challenge for LLM. Traditional position encoding techniques often struggle with long-range dependencies or varying-length sequences. To address this, methods like Contextual Position Encoding (CoPE) (Golovneva et al., 2024) dynamically adjust positional encodings based on context, improving performance on tasks requiring long-range dependencies. Training-free techniques, such as contextual sliding windows (An et al., 2024), extend the context length of pre-trained models without retraining, reducing computational overhead while maintaining sequence-processing capabilities. Parallel Context Windows (PCW) (Ratner et al., 2023) further enhances efficiency by splitting sequences into chunks for parallel processing. However, models tend to underutilize information from earlier inputs, favoring the middle or end (Liu et al., 2024b). Some methods like ThoT (Zhou et al., 2023) address this by improving reasoning across fragmented contexts, while LLMLingua (Jiang et al., 2023) compresses prompt representations, striking a balance between efficiency and context retention.
@@ 3 Background and Notation
The visual dependency ${\bm{D}}_{\text{vis}}(k)$ measures the extent to which the language model relies on visual features ${\bm{V}}_{I}$ when generating the $k$-th token. Specifically, the attention weights $\alpha_{k,i}$ at position $k$ are distributed among all available inputs, i.e., textual embeddings ${\bm{H}}_{T}$ and visual features ${\bm{V}}_{I}$. As the context length $m$ increases, the attention weights are normalized over a larger set of tokens, resulting in smaller individual attention weights for the visual features, i.e.,
@@ 4 Analysis on Visual Dependency of LVLMs in Long-Context Reasoning
To analyze the visual dependency of LVLMs in long-context reasoning, we construct a long-context dataset based on the SVIT dataset (Zhao et al., 2023). SVIT dataset comprises images with multiple types of annotations, including conversation, complex reasoning, referring QAs, and detail description. These annotations can be concatenated to form long text annotations for images. We retain a total of 15,295 samples through random sampling and filtering. The samples are proportionally distributed into groups of 0.5k, 1k, 1.5k, 2k, and 2.5k based on text length. Each sample is processed in a question-answer format, with filtering to ensure that the objects in the answers do not appear in the context provided for the questions. For the analysis, we used the following LVLMs with 7B parameters: DeepSeek-VL (Lu et al., 2024), LLaVA (Liu et al., 2023), LLaVA-Next (Liu et al., 2024a), ViP-LLaVA (Cai et al., 2024), LLaVA-Next-Video (Zhang et al., 2024), and Video-LLaMA2 (Cheng et al., 2024).
@@ 4.2 Does Context Length Affect Visual Dependency?
Figure 1: Performance of various LVLMs on long-context reasoning.
To investigate whether context length affects the visual dependency of LVLMs, we conducted experiments on various LVLMs with progressively long contexts. As shown in Figure 1, we observe that the performance of various LVLMs declines as the context length increases. DeepSeek-VL experiences an approximate 28% drop in performance. In contrast, LVLMs trained on video data, i.e., LLaVA-NeXT-Video and Video-LLaMA2, exhibit greater robustness on long text, with relatively smaller reductions in performance. However, their performance also diminishes with increased context length. It indicates that as the text length increases, the dependency of LVLMs on visual input decreases, leading to a decline in performance.
@@
Figure 2: Performance of various LVLMs on text-only long contexts.
@@ 4.3 Consistency Between Vision and Language Priors
To investigate the prediction consistency between vision and language priors, we analyze the disjoint portions of predictions with and without visual priors. Figure 3 presents the average performance of six models on different context lengths, comparing the accuracy of models provided visual priors (VL correct) and those only using language priors (Text correct). The proportion of samples that were correctly answered only when provided with an image generally decreases as the context length increases. It indicates that while visual dependency is more pronounced in shorter contexts, its impact diminishes as the model is exposed to more textual information. This suggests that the model increasingly relies on language reasoning under longer contexts. The proportion of samples only correctly answered without an image remains near zero across all context lengths. The results show that samples correctly answered using only language information remained to be correct even after adding visual information. It suggests that the models have effectively leveraged language priors to make accurate predictions, regardless of the availability of visual cues. As the context length increases, models rely more on the rich semantic information contained in the text, reducing visual dependency.
@@ 5 Context Pruning to Improve Visual Dependency
Figure 6: Overview of our method. Left: the token pruning process across multiple iterations; Right: pruning tokens in each LLM layer.
@@ 5.1 Context Pruning
where $Q^{(l)}_{i}$ and $K^{(l)}_{j}$ are the query and key representations of tokens $t_{i}$ and $t_{j}$, respectively, in layer $l$, and the attention weight $A^{(l)}_{ij}$ measures how much token $t_{i}$ attends to token $t_{j}$ in that layer. $\frac{1}{\sqrt{d_{k}}}$ is scaling factor, and $d_{k}$ the dimensionality of the key representations.
@@ 5.2 Theoretical Demonstration
where smaller entropy indicates higher attention concentration, meaning the model focuses more on specific tokens. Consider pruning the lowest $n\%$ of tokens based on attention weights. Let the remaining tokens form the pruned sequence $\bar{X}=[\bar{x}_{1},\bar{x}_{2},\cdots,\bar{x}_{\bar{T}}]$, where $\bar{T}=(1-n\%)\times T$. After pruning, the attention distribution of token $x_{i}$ is updated to $\bar{A}^{(l)}_{i}$, and the entropy after pruning are defined as:
@@
Since the low-weight tokens have been pruned, the remaining attention weights are concentrated on more important tokens, leading to a reduction in entropy:
@@
Thus, pruning reduces the entropy of attention distributions, making the model’s focus more stable and accurate. Furthermore, to quantify the attention stability, we can calculate the variance of the attention weights before and after pruning for token $x_{i}$, i.e.,
@@
where $\sigma^{2}(A^{(l)}_{i})$ and $\sigma^{2}(\bar{A}^{(l)}_{i})$ are the variance of the attention weights before and after pruning, respectively. Since pruning low-weight tokens increases the concentration of attention on the remaining tokens, the variance of the attention distribution increases:
@@
where the increase in variance indicates that the remaining attention is more focused on important tokens.
@@
Consider a LVLM processes both visual input $V$ and textual input $T$ to produce an output $O$, such that $O=f(V,T)$. The model’s dependence on each modality can be quantified through its attention mechanism and the sensitivity of the output to each input. Let $A_{V}$ and $A_{T}$ represent the attention distributions over visual and textual inputs, respectively, satisfying the normalization condition:
@@
When textual tokens are pruned, the textual input reduces from $T$ to a pruned version $\bar{T}$, containing fewer tokens. Consequently, the attention weights over textual inputs decrease, leading to a reduced $\beta$:
@@
where $\alpha^{\prime}>\alpha$ indicates increased dependence on visual inputs. This increase in $\alpha$ directly translates to greater sensitivity of the output $O$ to the visual input $V$:
@@
Similarly, the sensitivity to the textual input decreases:
@@
Therefore, as textual tokens are pruned, the model reallocates its attention from textual to visual inputs, resulting in increased $\alpha^{\prime}$ and decreased $\beta^{\prime}$. This shift enhances the model’s dependency on visual information for reasoning, as evidenced by both the attention weights and the sensitivity measures.

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) excel in cross-model tasks but experience performance declines in long-context reasoning due to overreliance on textual information and reduced visual dependency. In this study, we empirically analyze LVLMs in long-context reasoning, revealing that increased context length leads to a higher dependence on language at the expense of visual dependency. To address this issue, we propose a novel training-free context pruning method that selectively removes less critical textual information. Our approach enhances visual dependency and reduces textual noise, thereby improving LVLM performance in long-context reasoning. We validate our method by constructing a long-context dataset, demonstrating its effectiveness across various LVLMs. Moreover, we compare our method with token-pruning methods, demonstrating superior performance, and further analysis confirms the robustness of our method.
@@
Yucheng Zhou1      Zhi Rao2      Jun Wan2      Jianbing Shen1thanks:  Corresponding author. 1 SKL-IOTSC, CIS, University of Macau 2 Macau University of Science and Technology yucheng.zhou@connect.um.edu.mo, jianbingshen@um.edu.mo
@@ 1 Introduction
Recently, Large Language Models (LLMs) have significantly enhanced their performance in long-context reasoning through contextual position encoding and optimized attention strategies (Ding et al., 2024; Jin et al., 2024). These methods enable LLMs to retain and effectively utilize more information over long-context interactions, thereby maintaining contextual coherence and improving response accuracy. In contrast, LVLMs struggle to leverage visual information appropriately when processing long contexts. Therefore, LVLMs are required to integrate visual and textual information under long-context reasoning accurately.
@@
To improve the visual dependency in long-context reasoning, we propose a novel training-free context pruning method for LVLMs, which strategically reduces the interference of less critical textual information while preserving visual information. Our approach leverages the model’s internal attention mechanisms to identify and eliminate the trivial textual tokens based on their aggregated attention weights across multiple attention heads. By systematically pruning these less significant tokens, our method effectively diminishes textual noise and encourages more attention to visual information, enhancing visual dependency under long-context reasoning. In addition, we provide a theoretical foundation demonstrating that context pruning leads to more concentrated and stable attention distributions, which in turn fosters a stronger visual dependency on visual inputs. This theoretical demonstration is further validated through a detailed analysis of information flow.
To comprehensively evaluate the performance of LVLMs in long-context reasoning, we construct a long-context dataset based on the SVIT dataset (Zhao et al., 2023). Experimental results demonstrate the effectiveness of our approach across various LVLMs. Moreover, we compare our method with other token pruning approaches on the MMDU dataset, achieving superior performance. Our method demonstrates significant improvements across different model sizes. Additionally, we further verify the robustness of our method by comparing different token pruning strategies, as well as various pruning layers and rates. The characteristics of our pruning method are further analyzed through part-of-speech observations of the pruned tokens. Lastly, we conduct an exploration of the scaling laws about the relationship between token pruning rates and context length.
@@
  • We perform theoretical and information flow analysis demonstrating context pruning leads to more concentrated and stable attention distributions, which enhances visual dependency.
@@ 2 Background and Notation
The visual dependency ${\bm{D}}_{\text{vis}}(k)$ measures the extent to which the language model relies on visual features ${\bm{V}}_{I}$ when generating the $k$-th token. Specifically, the attention weights $\alpha_{k,i}$ at position $k$ are distributed among all available inputs, i.e., textual embeddings ${\bm{H}}_{T}$ and visual features ${\bm{V}}_{I}$. As the context length $m$ increases, the attention weights are normalized over a larger set of tokens, resulting in smaller visual attention weights, i.e.,
@@
Related Work (Summarized).
Recent advancements in large vision-language models (LVLMs) focus on enhancing cross-modal understanding by integrating visual and textual inputs for complex tasks (Wang et al., 2023b; Chen et al., 2023a, b). Models like LLaVA (Liu et al., 2023) and LLaVA-NeXT (Liu et al., 2024a) improve performance in multimodal reasoning and OCR, while approaches like DeepSeek-VL (Lu et al., 2024) address challenges of noisy inputs. In long-context reasoning, techniques like Contextual Position Encoding (CoPE) (Golovneva et al., 2024) and sliding windows (An et al., 2024) extend context handling, with methods such as ThoT (Zhou et al., 2023) and LLMLingua (Jiang et al., 2023) improving reasoning across fragmented contexts. The full version can be found in Appendix A.
@@ 3 Analysis on Visual Dependency of LVLMs in Long-Context Reasoning
To analyze the visual dependency of LVLMs in long-context reasoning, we construct a long-context dataset [^1] The dataset and code will be made publicly available after the paper’s acceptance to encourage community adoption. based on the SVIT dataset (Zhao et al., 2023). SVIT dataset comprises images with multiple types of annotations, including conversation, complex reasoning, referring QAs, and detail description. These annotations can be concatenated to form long text annotations for images. We retain a total of 15,295 samples through random sampling and filtering. The samples are proportionally distributed into groups of 0.5k, 1k, 1.5k, 2k, and 2.5k based on text length. Each sample is processed in a question-answer format, with filtering to ensure that the objects in the answers do not appear in the context provided for the questions. For the analysis, we used the following LVLMs with 7B parameters: DeepSeek-VL (Lu et al., 2024), LLaVA (Liu et al., 2023), LLaVA-Next (Liu et al., 2024a), ViP-LLaVA (Cai et al., 2024), LLaVA-Next-Video (Zhang et al., 2024), and Video-LLaMA2 (Cheng et al., 2024). The performance metrics on this dataset refer to accuracy, measured by the correctness of inference results.
@@ 3.2 Does Context Length Affect Visual Dependency?
Figure 1: Performance of LVLMs on long-context reasoning.
To investigate whether context length affects the visual dependency of LVLMs, we conducted experiments on various LVLMs with progressively long contexts. As shown in Figure 1, we observe that the performance of various LVLMs declines as the context length increases. DeepSeek-VL experiences an approximate 28% drop in performance. Moreover, LVLMs trained on video data, i.e., LLaVA-NeXT-Video and Video-LLaMA2, exhibit greater robustness on long text, with relatively smaller reductions in performance. It indicates as the text length increases, the dependency of LVLMs on visual input decreases, leading to a decline in performance.
@@
Figure 2: Performance of LVLMs on text-only long contexts.
@@ 3.3 Consistency Between Vision and Language Priors
We analyze the consistency between vision and language priors by comparing prediction performance with and without visual input across varying context lengths. Figure 3 illustrates the average accuracy of six models, showing that samples correctly answered with language priors alone remained accurate even when visual cues were added, highlighting the robustness of language reasoning. The detailed analysis can be found in the Appendix B.
@@ 4 Context Pruning to Improve Visual Dependency
Figure 6: Overview of our method. Left: token pruning across multiple iterations; Right: pruning tokens in each LLM layer.
@@ 4.1 Context Pruning
where $Q^{(l)}_{i}$ and $K^{(l)}_{j}$ are the query and key representations of tokens $t_{i}$ and $t_{j}$, respectively, in layer $l$, and the attention weight $A^{(l)}_{ij}$ measures how much token $t_{i}$ attends to token $t_{j}$ in that layer. $\frac{1}{\sqrt{d_{k}}}$ is scaling factor, and $d_{k}$ the dimensionality of $K^{(l)}_{j}$.
@@ 4.2 Theoretical Demonstration
where smaller entropy indicates higher attention concentration, meaning the model focuses more on specific tokens. Consider pruning the lowest $n\%$ of tokens based on attention weights. Let the remaining tokens form the pruned sequence $\bar{X}=[\bar{x}_{1},\bar{x}_{2},\cdots,\bar{x}_{\bar{T}}]$, where $\bar{T}=(1-n\%)\times T$. After pruning, the attention distribution of token $x_{i}$ is updated to $\bar{A}^{(l)}_{i}$, and the entropy are defined as:
@@
Since the low-weight tokens have been pruned, remaining weights are concentrated on more important tokens, leading to entropy reduction:
@@
Thus, pruning reduces the entropy of attention distributions, making the model’s focus more stable and accurate. To quantify the attention stability, we can calculate the variance of the attention weights before and after pruning for token $x_{i}$, i.e.,
@@
where $\sigma^{2}(A^{(l)}_{i})$ and $\sigma^{2}(\bar{A}^{(l)}_{i})$ are the variance of the attention weights before and after pruning, respectively. Since pruning low-weight tokens increases the concentration of attention on the remaining tokens, variance of attention distribution increases:
@@
where the increase in variance indicates the remaining attention is more focused on important tokens.
@@
Consider a LVLM processes both visual input $V$ and textual input $T$ to produce an output $O$, such that $O=f(V,T)$. The model’s dependence on each modality can be quantified through its attention mechanism and the sensitivity of the output to each input. Let $A_{V}$ and $A_{T}$ represent the attention distributions over visual and textual inputs, satisfying the normalization condition:
@@
When textual tokens are pruned, the textual input reduces from $T$ to a pruned version $\bar{T}$, containing fewer tokens. The attention weights over textual inputs decrease, leading to a reduced $\beta$:
@@
where $\alpha^{\prime}>\alpha$ indicates increased dependence on visual inputs. This increase in $\alpha$ directly translates to greater sensitivity of output $O$ to visual input $V$:
@@
Similarly, sensitivity to textual input decreases:
@@
Therefore, as textual tokens are pruned, the model reallocates its attention from textual to visual inputs, resulting in increased $\alpha^{\prime}$ and decreased $\beta^{\prime}$. This shift enhances the model’s dependency on visual information for reasoning, as evidenced by both the attention weights and sensitivity measures.
4.3 Information Flow Analysis
Following the approaches Zhou et al. (2024); Wang et al. (2023a), we further validate the effectiveness of our method through information flow. The saliency score for each element in the attention matrix is computed using an adapted version of the Taylor expansion method Michel et al. (2019), i.e.,
$I_{l}=\sum_{h}\left|{\bm{A}}_{h,l}^{\top}\odot\frac{\partial{\mathcal{L}}(x)}{\partial{\bm{A}}_{h,l}}\right|,$ (17)
where $h$ and $l$ denote distinct attention heads and Transformer layers, respectively, and ${\mathcal{L}}(x)$ represents the loss function. We propose six significance scores based on information flow, i.e.,
$S_{vr}$ $=\frac{\sum_{(i,j)\in C_{vr}}I_{l}(i,j)}{|\{(v,r_{k}):k\in[1,R]\}|},$ (18)
$S_{vc}$ $=\frac{\sum_{(i,j)\in C_{vc}}I_{l}(i,j)}{|\{(v,c_{k}):k\in[1,C]\}|},$ (19)
$S_{rt}$ $=\frac{\sum_{(i,j)\in C_{rt}}I_{l}(i,j)}{|\{(r_{k},t):k\in[1,R]\}|},$ (20)
$S_{ct}$ $=\frac{\sum_{(i,j)\in C_{ct}}I_{l}(i,j)}{|\{(c_{k},t):k\in[1,C]\}|},$ (21)
$S_{vt}$ $=\frac{\sum_{(i,j)\in C_{vt}}I_{l}(i,j)}{|\{(v,t)\}|},$ (22)
$S_{ww}$ $=\frac{\sum_{(i,j)\in C_{ww}}I_{l}(i,j)}{|C_{ww}|},$ (23)
where $v,r_{k},c_{k},t$, $R$, and $C$ represent the visual part, preserved tokens, pruned tokens, the target position, the total number of preserved tokens, and the total number of pruned tokens, respectively.
Model Method Context Length (token)
0.5k 1.0k 1.5k 2.0k 2.5k
DeepSeek-VL (Lu et al., 2024) Baseline 75.60 66.20 49.00 43.90 47.80
Ours 78.85(3.25) 71.98(5.78) 67.87(18.87) 66.33(22.43) 62.63(14.83)
ViP-LLaVA (Cai et al., 2024) Baseline 84.40 79.80 70.10 66.70 63.60
Ours 91.57(7.17) 91.50(11.70) 87.28(17.18) 83.79(17.09) 75.74(12.14)
LLaVA (Liu et al., 2023) Baseline 91.10 88.90 81.80 77.30 74.40
Ours 95.91(4.81) 95.36(6.46) 93.92(12.12) 91.40(14.10) 87.05(12.65)
LLaVA-Next (Liu et al., 2024a) Baseline 97.50 95.10 90.90 90.30 85.60
Ours 97.61(0.11) 95.62(0.52) 91.47(0.57) 90.55(0.25) 86.60(1.00)
Video-LLaMA2 (Cheng et al., 2024) Baseline 95.60 95.10 92.00 90.70 90.00
Ours 95.78(0.18) 97.42(2.32) 97.12(5.12) 96.70(6.00) 96.31(6.31)
LLaVA-Next-Video (Zhang et al., 2024) Baseline 95.90 95.20 95.10 95.00 93.80
Ours 96.08(0.18) 95.52(0.32) 95.59(0.49) 95.13(0.13) 94.12(0.32)
Table 1: Performance comparison of our method and baselines across various LVLMs. The blue numbers in parentheses indicate the performance gain achieved by our method compared to the baselines.
Figure 7: Information flow results.
Method LLaVA-Next-Video Video-LLaMA2 LLaVA-Next
Baseline 44.0 43.7 43.9
Ours 46.1(2.10) 45.5(1.80) 44.3(0.40)
Table 2: Performance comparison (Avg.) between baselines and our method on MMDULiu et al. (2024c).
Figure 8: Performance (Avg.) across different sizes of LLaVA-Next on MMDU.
Method Avg.
Baseline 43.9
SpAtten 43.5
LTP 44.0
Ours 44.3
Table 3: Performance comparison of our method, SpAtten Wang et al. (2021), and LTP Bonnaerens and Dambre (2023) on MMDU with LLaVA-Next.
The set $C_{ww}$ encompasses all token pairs except those already included in $C_{vr}$, $C_{vc}$, $C_{rt}$, $C_{ct}$, or $C_{vt}$. Specifically, $S_{vr}$ represents the flow from the visual part to preserved tokens, and $S_{vc}$ to pruned tokens. $S_{rt}$ and $S_{ct}$ capture the flows from preserved and pruned tokens to the target position, respectively. $S_{vt}$ denotes the direct flow from the visual part to the target, while $S_{ww}$ quantifies flow among all tokens, excluding those covered by the above flow.
Figure 7 shows that pruning enhances the flow from visual inputs to preserved tokens ($S_{vr}$) and the target ($S_{vt}$), while reducing flow to pruned tokens ($S_{vc}$). These changes validate Proposition 1 by demonstrating that pruning stabilizes and concentrates attention on important tokens, reducing entropy. They also support Proposition 2, as the increased $S_{vr}$ and $S_{vt}$ highlight the model’s shift towards greater dependency on visual inputs. More Analysis can be found in Appendix C.